OpenStudy (anonymous):

which of the following correctly describes the domain of the function shown? r(x)=2x/x^2-1 (x cannot equal +-1, x cannot equal 0, all real numbers, or x cannot equal )

OpenStudy (anonymous):

a zero in the denominator makes the function undefined. we need to set the denominator equal to zero to find what x shouldnt be..... 0=x^2-1 1=x^2 sqrt1=x x cannot be +-1

OpenStudy (anonymous):

oh I see thanks :)